# T-MAE: Temporal Masked Autoencoders for Point Cloud Representation Learning

### v1

@@ Abstract
The scarcity of annotated data in outdoor point cloud segmentation poses a significant obstacle in harnessing the modeling capabilities of advanced networks like transformers. Consequently, scholars have been actively investigating efficacious self-supervised pre-training strategies, e.g. contrasting learning and reconstruction-based pretext tasks. Nevertheless, temporal information, which is inherent in the LiDAR point cloud sequence, is consistently disregarded. To better utilize this property, we propose an effective pre-training strategy, namely Temporal Masked Auto-Encoders (T-MAE), which takes as input temporally adjacent frames and learns temporal dependency. A SiamWCA backbone, containing a Siamese encoder and a window-based cross-attention (WCA) module, is established for the two-frame input. Taking into account that the motion of an ego-vehicle alters the illumination angles of the same instance, temporal modeling also serves as a robust and natural data augmentation, enhancing the comprehension of target objects. Moreover, instead of utilizing consecutive frames, it is more cost-effective and powerful by using distant historical frames. SiamWCA is a powerful architecture but heavily relies on annotated data. With our T-MAE pre-training strategy, we achieve the best performance on the Waymo dataset among self-supervised learning methods. Comprehensive experiments are conducted to validate all components of our proposal. Upon acceptance, the source code will be made accessible.
@@ 1 Introduction
Figure 1: T-MAE performance on Waymo dataset [39]. Our T-MAE pre-training achieves substantially better results than random initialization. It also achieve better results than the SOTA SSL method MV-JAR [48] with significantly fewer iterations. Each point triplet corresponds to a specific fine-tuned model variant and shows the performance difference with two arrows. From left to right, the triplets correspond to a model fine-tuned with 5%, 10% and 20% data.
@@
In the domains of language and vision, researchers have observed that as models become more complex, the available amount of labeled data is often insufficient to adequately train huge models [11, 18] e.g. Vision Transformer (ViT) [12]. Consequently, there is a growing interest in exploring self-supervised learning (SSL) approaches as a potential solution to overcome this limitation. SSL serves as a pre-training technique with unlabeled data, accelerating the convergence of the models and improving their performance for downstream tasks [16, 6, 9]. The same challenge extends to the domain of point clouds, where annotations are more costly and time-consuming to obtain [46, 14]. For instance, a mere 10% and 0.8% of sweeps are annotated in the nuScene [2] and ONCE [29] datasets. This challenge makes pre-training for point cloud understanding a non-trivial endeavor.
@@
Prior works mainly focus on synthetic and isolated objects [50, 56, 54, 35, 27] and indoor scene understanding [19, 24, 49, 36]. Transferring these methods to outdoor LiDAR points is challenging due to their sparsity and dynamic environmental conditions. At present, most self-supervised methods used for understanding point clouds in autonomous driving rely on contrastive learning [47, 53, 32, 34, 33]. These approaches model the similarity and dissimilarity between entities, such as segments [33, 45] and/or points [47]. In the wake of masked image modeling as a pretext task [18], efforts have also been devoted to the reconstruction of masked points [31, 41, 52, 48]. The main idea is randomly masking points or voxels and urging the network to infer the coordinates of points [52] and/or voxels [48] or other properties, e.g. occupancy [31, 1] and curvature [41]. Nevertheless, these methodologies often operate within the confines of a single-frame scenario, disregarding the fact that LiDAR data is typically acquired on a frame-by-frame basis. In other words, the valuable semantic information that is introduced by temporally adjacent frames is often not exploited.
@@
Several methods attempt to leverage temporal information [21, 45, 33, 25] by incorporating multi-frame input during the self-supervised phase but their core concepts remain grounded in contrastive learning. Specifically, the scans captured at different times are treated as augmented samples of the same scene, without including temporal correspondence into the modeling procedure.
@@
Therefore, we propose a new self-supervised method, namely T-MAE, to exploit the accumulated observations. The key idea of T-MAE is to learn a powerful representation for point cloud scans and integrate features from the past for present use. During the pre-training stage, the current scan is voxelized with a high masking ratio, while the previous scan is fed entirely to the encoder. Then, the reconstruction pretext task is performed by incorporating three types of information: voxel embeddings of the past scan, visible voxel embeddings of the current scan, and the position of masked voxels. In this way, the WCA module learns to incorporate historical information into the current frame using unlabeled data. The proposed T-MAE pre-training strategy endows the network with both a powerful representation for sparse point clouds and the capacity to strengthen the present by learning from the past. As shown in Fig. 1, the proposed T-MAE achieves higher overall and pedestrian-specific mAPH than the randomly initialized baseline, namely the same model but trained from scratch. Moreover, T-MAE also outperforms state-of-the-art MV-JAR [48] with over 1.6×1.6$\times$1.6 × fewer training iterations. Our contributions are summarized as follows:
  • We propose T-MAE, a novel and effective SSL approach for sparse point clouds, that learns temporal modeling in the process of reconstructing masked points.
  • We design a SiamWCA backbone, containing a Siamese encoder and a window-based sparse cross-attention (WCA) module, to incorporate historical information.
  • Rather than concatenating multiple consecutive point cloud scans, our approach demonstrates that incorporating more distant historical frames provides greater advantages.
  • Our experiments demonstrate the efficacy of T-MAE by attaining impressive results on the Waymo Open dataset. T-MAE achieves 55.71 mAPH for pedestrians with only 5% of the labeled data and surpasses the SOTA SSL approaches by a substantial margin across all levels of fine-tuning data amount.
@@ 2.1 Static Self-supervised Learning
SSL for point clouds is a burgeoning field due to the scarcity of annotations. In general, the pipeline of SSL consists of two phases. During the initial stage, namely pre-training with unlabeled data, the network is trained using a pretext task. Subsequently, in downstream tasks such as segmentation and detection, the pre-trained weights are loaded to the backbone, and the backbone is attached with task-specific heads. The resulting model is further fine-tuned using annotated data.
@@
The initial research efforts are primarily directed towards the exploration of contrastive learning. The underlying hypothesis is that an image or 3D scene demonstrates feature equivalence even after undergoing different transformations [9, 5]. One of the key challenges in the point cloud domain is establishing correspondences across scenes. PointContrast [47] and DepthContrast [57] track the points while performing different transformations. GCC-3D [25] exploit sequential information to obtain pseudo instances and then perform contrasting on the instance level. SegContrast [32] firstly obtains segments by an unsupervised clustering and then contrasting segments between two transformed views. These methods allow the network to learn equivalence with regard to geometric transformations. However, contrastive learning-based SSL approaches usually suffer from careful tuning of hyperparameters and complicated pre- or post-processing to find correspondences.
@@
The focus has shifted to reconstructing masked points, following the success of masked image modeling [18, 44]. The masking strategies remain consistent, i.e. voxelize a point cloud into cubes followed by random masking of these cubes. The reconstruction targets vary across papers. OccupancyMAE [31] classifies voxels to whether they are occupied. Geo-MAE [41] infers occupancy as well as the normal and the curvature of each masked voxel. The GD-MAE’s pretext task is to reconstruct a fixed number of points for masked voxels [52]. MV-JAR [48] partitions the masked Voxel into two categories, necessitating the network to either predict the coordinates of the voxels or produce the points themselves. In this paper, we employ the same masking strategy as GD-MAE [52] but expand the knowledge source to include both the current visible voxels and voxels from a previous frame as a reference.
2.2 Temporal Self-supervised Learning
@@
There are many SSL methods designed to address temporal or spatiotemporal tasks. In video-related tasks, such as MAE-ST [15], a random patch masking strategy is employed over consecutive frames. The model is tasked with recovering these masked patches while considering information from neighboring frames. VideoMAE [42] retains the same reconstruction pipeline as the pretext task but exploits a tube masking strategy. SiameseMAE [17] learns object-centered representations with the help of cross-attention layers and an asymmetrical masking technique on consecutive frames. There are also content-based masking strategies, e.g. motion-guided masking [20, 30]. The primary concept underlying these methods is acquiring an understanding of temporal dependency across frames through the process of reconstructing patches with reference to consecutive frames.
@@
While this concept is effectively used in video understanding, it is rarely applied to the learning of sparse point cloud representations. This is because point clouds are typically handled on a frame-by-frame basis rather than being regarded as a temporal sequence. For instance, when using temporally adjacent scans as input during self-supervised learning, both STSSL [45] and TARL [33] use HDBSCAN [3] to obtain segments and apply self-supervision by minimizing the feature distance between the same segments or points between neighbor frames. While they achieve impressive results, two issues persist 1) The use of elaborate and time-consuming pre-processing methods to obtain segments. 2) A primary focus on object consistency across frames rather than understanding object motion.
@@
Unlike other approaches, this paper aims to reduce the reliance on complex preprocessing techniques and emphasizes the significance of enabling the model to establish temporal correspondence through self-supervised learning from unlabeled data.
@@ 3 Method
Figure 2: Overview of the proposed T-MAE pre-training. Two frames are sampled from a sequence of point clouds according to temporal batch-based strategy. The prior scan $\mathcal{P}^{t+t_{1}}$ is voxelized and directly fed into the Siamese windowed cross-attention (SiamWCA) backbone. The current scan $\mathcal{P}^{t+t_{2}}$ undergoes a masking process before fed into the backbone. Furthermore, pillar-wise tokens are computed by a Siamese hierarchical encoder. The two-way gray arrow indicates weight sharing. The WCA module takes as input the full tokens of the previous frame and the partial observation of the current frame and creates enhanced tokens. In the diffusion process, masked tokens are updated by their neighbor tokens from WCA by a group of convolution layers. Subsequently, a reconstruction head recovers masked points. During fine-tuning, the full SiamWCA backbone is initialized with pre-trained weights obtained from the pre-training phase.
In this section, previous works are briefly discussed in Sec. 3.1. Then, the proposed T-MAE pre-training strategy is demonstrated in Sec. 3.2. Eventually, the window-based sparse cross-attention (WCA) module is elaborated in Sec. 3.3.
@@ 3.1 Preliminaries
Pillar-based representation and sparse regional self-attention.
@@
The pillar-based representation is an efficient representation introduced in PointPillars [23]. It divides 3D points into infinite-height voxels and computes pillar-wise features. These features can be treated as a pseudo-image in a bird’s eye view. On top of the pillar-based representation, SST [13] proposes a sparse regional self-attention (SRA) module to address a challenge in applying ViT to sparse LiDAR points. SST divides the pillars into non-overlapping windows with a fixed size and then applies self-attention to the non-empty pillars within the same window. The receptive field of each pillar is expanded through a window shift operation. T-MAE and many recent works [52, 48, 40] are built up on the basic block SRA.
Temporal batch.
@@
A sweep of a point cloud at time $t$ is denoted by $\mathcal{P}^{t}=\{(p^{t}_{k})\}_{k=1}^{K}$, where $p_{k}$ is a 3D point $p_{k}\in\mathbb{R}^{3}$. Moreover, a sequence of point clouds is an ordered set of $\mathcal{P}^{t}$, which is denoted by $\mathcal{P}=\{\mathcal{P}^{1},\mathcal{P}^{2},\ldots,\mathcal{P}^{T}\}$. During pre-training and fine-tuning, two scans are needed feed the Siamese encoder. Due to the high sampling frequency of the LiDAR system, two consecutive frames usually contain redundant duplicate information, which is verified in Sec. 5. Conversely, when there is a substantial time gap between two frames, their overlap may be limited, rendering the historical information less useful. Therefore, we follow TARL [33] to introduce the concept of a temporal batch so that the interval of two frames is constrained to an appropriate range. Specifically, we sample a batch of consecutive frames $\mathcal{B}^{t}=\{\mathcal{P}^{t+1},\mathcal{P}^{t+2},\ldots,\mathcal{P}^{t+n}\}$ from a sequence $\mathcal{P}$. Then, two frames, $\mathcal{P}^{t+t_{1}}$ and $\mathcal{P}^{t+t_{2}}$ are sampled from the first and last one-third of the batch. More specifically, $0\leq t_{1}<\frac{n}{3}$ and $\frac{2n}{3}\leq t_{2}$.

### v3

@@ Abstract
The scarcity of annotated data in LiDAR point cloud understanding hinders effective representation learning. Consequently, scholars have been actively investigating efficacious self-supervised pre-training paradigms. Nevertheless, temporal information, which is inherent in the LiDAR point cloud sequence, is consistently disregarded. To better utilize this property, we propose an effective pre-training strategy, namely Temporal Masked Auto-Encoders (T-MAE), which takes as input temporally adjacent frames and learns temporal dependency. A SiamWCA backbone, containing a Siamese encoder and a windowed cross-attention (WCA) module, is established for the two-frame input. Considering that the movement of an ego-vehicle alters the view of the same instance, temporal modeling also serves as a robust and natural data augmentation, enhancing the comprehension of target objects. SiamWCA is a powerful architecture but heavily relies on annotated data. Our T-MAE pre-training strategy alleviates its demand for annotated data. Comprehensive experiments demonstrate that T-MAE achieves the best performance on both Waymo and ONCE datasets among competitive self-supervised approaches. Codes will be released here.
Abstract
This supplementary provides more details and analysis of our method. The implementation details are illustrated in Sec. S1. Additional ablation results are provided in Sec. S2. Section S3 discusses the selection of temporal gap during inference. More quantitative comparisons with other methods on the Waymo dataset [49] are depicted in Sec. S4. The transferability of T-MAE is verified in Sec. S5 and a comparison with multi-frame non-SSL methods is provided in Sec. S6. We analyze what the WCA module learns from the T-MAE pre-training in Sec. S7. Sec. S8 provides details about the comparison of finetuning iterations. Eventually, more qualitative results and limitations are presented in Sec. S9 and Sec. S10.
Keywords:
Self-supervised learning LiDAR point cloud 3D detection
@@ 1 Introduction
[…]
Figure 1: T-MAE performance on Waymo [49]. Left: Each point triplet shows the performance differences to three models finetuned with the same data. The triplets show finetuned models with 8K, 16K, 32K labeled frames (left to right). Our T-MAE pre-training outperforms both random initialization and the SOTA SSL method MV-JAR [60] with significantly fewer iterations. Right: T-MAE yields higher mAPH for pedestrians when finetuned with half the labeled data than MV-JAR.
@@
As deep neural networks become more complex, the available amount of labeled data is often insufficient to adequately train huge models [14, 24], e.g., Vision Transformer (ViT) [15]. Consequently, there is a growing interest in exploring self-supervised learning (SSL) approaches as a potential solution to overcome this limitation. SSL serves as a pre-training technique with unlabeled data, accelerating the convergence of the models and improving their performance for downstream tasks [20, 8, 11]. The same challenge extends to the domain of point clouds, where annotations are more costly and time-consuming to obtain [58, 17]. For instance, a mere 10% and 0.8% of frames are annotated in the nuScene [4] and ONCE [37] datasets, respectively. This challenge makes pre-training for point cloud understanding a non-trivial endeavour.
@@
Prior works mainly focus on synthetic and isolated objects [62, 71, 69, 43, 34, 28, 72] and indoor scene understanding [25, 31, 61, 45, 35]. Transferring these methods to outdoor LiDAR points is challenging due to their sparsity and dynamic environmental conditions. At present, most of the self-supervised methods used for understanding point clouds in autonomous driving rely on contrastive learning [59, 67, 40, 42, 41]. These approaches model the similarity and dissimilarity between entities, such as segments [41, 57] and/or points [59]. In the wake of masked image modeling as a pretext task [24], efforts have also been devoted to the reconstruction of masked points [39, 53, 64, 60]. The main idea is randomly masking points or voxels and urging the network to infer the coordinates of points [64] and/or voxels [60] or other properties, e.g., occupancy [39, 2] and curvature [53]. Nevertheless, these methodologies often operate within the confines of a single-frame scenario, disregarding the fact that LiDAR data is typically acquired on a frame-by-frame basis. In other words, the valuable semantic information in temporally adjacent frames is barely exploited.
@@
Several methods attempt to leverage temporal information [27, 57, 41, 32] by incorporating multi-frame input during the self-supervised phase but their core concepts remain grounded in contrastive learning. Specifically, the point clouds captured at different times are treated as augmented samples of the same scene, without including temporal correspondence into the modeling procedure.
@@
Therefore, we propose a new self-supervised paradigm, namely T-MAE, to exploit the accumulated observations. During the pre-training stage, the current scan is voxelized with a high masking ratio, while the previous scan is fed entirely to the encoder. Then, the pretext task is to reconstruct the current scan by incorporating voxel embeddings of the past scan, visible voxel embeddings of the current scan, and the position of masked voxels. This way, the proposed windowed cross-attention module learns to incorporate historical information into the current frame using unlabeled data. The T-MAE pre-training strategy endows the network with both a powerful representation for sparse point clouds and the capacity to strengthen the present by learning from the past. As shown in Fig. 1, the proposed T-MAE achieves higher overall and pedestrian-specific mAPH than the randomly initialized baseline, namely the same model but trained from scratch. Moreover, T-MAE also outperforms state-of-the-art MV-JAR [60] with over 1.6×1.6$\times$1.6 × to 2.4×2.4$\times$2.4 × fewer finetuning iterations.
Our contributions are summarized as follows: 1) We propose T-MAE, a novel and effective SSL approach for representation learning of sparse point clouds, that learns temporal modeling in the process of reconstructing masked points. 2) We design a SiamWCA backbone, containing a Siamese encoder and a windowed sparse cross-attention (WCA) module, to incorporate historical information. 3) Our experiments demonstrate the efficacy of T-MAE by attaining substantial improvements on the Waymo and ONCE datasets. Notably, T-MAE with 5% labeled data outperforms the SOTA SSL approach MV-JAR in terms of mAPH for pedestrians, even if MV-JAR employs 10% labeled data.
@@ 2 Related Work
Static Self-supervised Learning. SSL for point clouds is a burgeoning field due to the scarcity of annotations. In general, the pipeline of SSL consists of two phases. First, the network is trained using a pretext task with unlabeled data. Second, in downstream tasks such as segmentation and detection, the pre-trained weights are loaded to the backbone, and the backbone is attached with task-specific heads. The resulting model is finetuned with annotated data.
@@
The initial research efforts are primarily directed towards contrastive learning. The underlying hypothesis is that an image or 3D scene demonstrates feature equivalence even after undergoing different transformations [11, 7]. One of the key challenges in the point cloud domain is establishing correspondences across scenes. PointContrast [59] and DepthContrast [74] track the points while performing different transformations. GCC-3D [32] exploit sequential information to obtain pseudo instances and then perform contrasting on the instance level. SegContrast [40] firstly obtains segments by an unsupervised clustering and then contrasts segments between two transformed views. These methods allow the network to learn equivalence regarding geometric transformations. However, contrastive SSL approaches usually suffer from careful tuning of hyperparameters and complicated pre- or post-processing to find correspondences.
@@
The focus has shifted to reconstructing masked points, following the success of masked image modeling [24, 56]. The masking strategies remain consistent, i.e. voxelize a point cloud into cubes followed by random masking of these cubes. The reconstruction targets vary across papers. OccupancyMAE [39] classifies voxels to whether they are occupied. Geo-MAE [53] infers occupancy as well as the normal and curvature of each masked voxel. The GD-MAE’s pretext task is to reconstruct a fixed number of points for masked voxels [64]. MV-JAR [60] partitions the masked Voxel into two categories, necessitating the network to either predict the coordinates of the voxels or produce the points themselves. In this paper, we employ the same masking strategy as GD-MAE [64] but expand the knowledge source to include both the current visible voxels and voxels from a previous frame as a reference.
@@
Temporal Self-supervised Learning. There are many SSL methods designed to address temporal or spatiotemporal tasks in the video domain. MAE-ST [18] employs a random patch masking strategy over consecutive frames. The model is tasked with recovering these masked patches while considering information from adjacent frames. VideoMAE [54] retains the same reconstruction pipeline as the pretext task but exploits a tube masking strategy. SiameseMAE [22] learns object-centred representations with the help of cross-attention layers and an asymmetrical masking technique on consecutive frames. There are also content-based masking strategies, e.g., motion-guided masking [26, 38]. The primary concept underlying these methods is acquiring an understanding of temporal dependency across frames through the process of reconstructing patches with reference to consecutive frames.
@@
While this concept is effectively used in video understanding, it is rarely applied to learn sparse point cloud representations. This is because point clouds are typically handled on a frame-by-frame basis rather than being regarded as a temporal sequence. For instance, when using temporally adjacent scans as input during self-supervised learning, both STSSL [57] and TARL [41] use HDBSCAN [5] to obtain segments and apply self-supervision by minimizing the feature distance between the same segments or points of neighbour frames. While they achieve impressive results, two issues persist 1) The use of elaborate and time-consuming pre-processing methods to obtain segments. 2) A primary focus on object consistency across frames rather than understanding object motion.
@@
Unlike other approaches, we aim to reduce the reliance on complex preprocessing techniques and focus on enabling the model to establish temporal correspondence through self-supervised learning from unlabeled data.
@@ 3 Method
We first briefly discuss important preliminaries in Sec. 3.1 from previous works although they mainly focus on the single-frame setting. To incorporate historical frames, we build up our framework as in Fig. 3. The key component of the framework, SiamWCA, is introduced in 3.2. Then, the proposed T-MAE pre-training strategy is described in Sec. 3.3 and the corresponding windowed sparse cross-attention (WCA) module is elaborated in Sec. 3.4.
@@ 3.1 Preliminaries
Pillar-based representation and sparse regional self-attention. The pillar-based representation is an efficient representation introduced in PointPillars [30]. It divides 3D points into infinite-height voxels and computes pillar-wise features. These features can be treated as a pseudo-image in a bird’s eye view. On top of the pillar-based representation, SST [16] proposes a sparse regional self-attention (SRA) module to address challenges in applying ViT to sparse LiDAR points. SST divides the pillars into non-overlapping windows and then applies self-attention within each window. The receptive field of each pillar is expanded through a window shift operation. This window partition and shift make the SRA module achieve a good balance of efficiency and accuracy. T-MAE and many recent works [64, 60, 50] are built upon the basic SRA block.
Reconstruction pretext task and single-frame baseline. The original MAE [24] randomly masks patches of an image and employs a ViT to reconstruct the image. To perform this reconstruction pretext task in the LiDAR domain, we follow the concept of the state-of-the-art SSL method GD-MAE [64] which operates on a single frame. The framework encompasses several key stages, i.e. voxelization, masking, encoding, dense feature recovery, and reconstruction. Through this pretext task, GD-MAE acquires valuable weights for the encoder, which are later utilized for downstream tasks.
$t$ ${t-3}\ \sim\ t$ $t$ ${t-3}\ \sim\ t$
Refer to caption Refer to caption Refer to caption Refer to caption
(a) A stationary vehicle. (b) A moving vehicle.
Figure 2: Comparison between single- and four-frame concatenation. While simple frame concatenation generally improves point density and detection rates, it can introduce spurious points in non-static scene parts that may degrade the detection performance. Since we combine consecutive frames via learned cross-attention, our approach is less affected by this problem. The blue bounding boxes indicate the ground truth for the current frame.
Analysis. Up to this point, a common approach has been to employ a single-frame reconstruction pretext task for LiDAR points. However, incorporating historical frames poses a non-trivial challenge. One straightforward approach is to concatenate two point clouds after aligning them with ego-poses, similar to the three-frame variant of SST [16]. However, as shown in Fig. 2, while concatenation helps to identify static objects effectively, it introduces challenges for detectors when dealing with moving objects. We provide quantitative results in Tab. 3. To address this problem, we suggest learning from past data in a latent space.
3.2 Framework Overview
Figure 3: Overview of our architecture and the proposed T-MAE pre-training. Two frames are sampled from a sequence of point clouds and are voxelized. During pre-training, the current frame $\mathcal{P}^{t_{2}}$ undergoes an additional masking process. Note that the dashed boxes indicate operations for pre-training phase only. Next, voxel-wise tokens are computed by a Siamese encoder. The two-way gray arrow indicates weight sharing. The WCA module takes as input the full tokens of the previous frame and the partial observation of the current frame and outputs enhanced tokens. The dense feature recovery places sparse tokens back to a dense feature map and convolves the map to fill empty locations. Subsequently, the feature map is either fed to a reconstruction head that recovers masked points, or to a detection head predicting bounding boxes.
As illustrated in Fig. 3, the proposed framework compromises voxelization, encoding, windowed cross-attention, dense feature recovery and two separate heads for pre-training and detection, respectively. This subsection elaborates on the key components for supervised learning, e.g., training from scratch and fine-tuning, whereas SSL-related components are introduced in Sec. 3.3.
@@ 3.2 Framework Overview
Temporal batch-based sampling. Consider a sequence of point clouds as an ordered set of point clouds, denoted as $\mathbf{P}=\{\mathcal{P}^{1},\mathcal{P}^{2},\ldots,\mathcal{P}^{t},\ldots,\mathcal{P}^{T}\}$. In this ordered set, $\mathcal{P}^{t}=\{(p^{t}_{k})\}_{k=1}^{K}$ is a sweep of point cloud at time $t$, where $p_{k}$ is a 3D point $p_{k}\in\mathbb{R}^{3}$ and $K$ indicates the number of points. In our setting, two frames are needed to feed the network. Due to the high sampling frequency of the LiDAR system, two consecutive frames usually contain redundant duplicate information, which is verified in Sec. S3. Conversely, when there is a substantial time gap between two frames, their overlap may be limited, making the historical information less useful. Therefore, we follow TARL [41] to introduce the concept of a temporal batch so that the interval of two frames is constrained to an appropriate range. Specifically, we sample a batch of consecutive frames $\mathcal{B}^{t}=\{\mathcal{P}^{t+1},\mathcal{P}^{t+2},\ldots,\mathcal{P}^{t+n}\}$ from a sequence $\mathcal{P}$. Then, two frames, $\mathcal{P}^{t_{1}}$ and $\mathcal{P}^{t_{2}}$ are sampled from the first and last one-third of the batch, namely, $t_{1}\in\{t+1,t+2,\ldots,\left\lfloor t+\frac{n}{3}\right\rfloor\}$ and $t_{2}\in\{\left\lceil t+\frac{2n+1}{3}\right\rceil,\ldots,n-1,n\}$.
Alignment and voxelization. Given two frames, the previous frame $\mathcal{P}^{t_{1}}$ is transformed to the coordinate system of the current frame $\mathcal{P}^{t_{2}}$ by their ego-poses. The pose information is available in most datasets [49, 4, 37] or easily obtained by means of GPS/IMU, odometry approaches [9], structure from motion (SfM) algorithms [46, 47], or SLAM systems [13]. Next, each point cloud is divided into discrete voxels. Two linear layers map point coordinates to high-dimensional features and a voxel-wise representation is obtained via voxel-wise average pooling. The voxels can be regarded as pillars, owing to the infinite height.
A Siamese encoder and windowed cross-attention (SiamWCA). A Siamese encoder [3] is a two-branch network where both branches share the same configurations and weights. In this work, it is utilized to encode the pillar-wise representations of both frames to sparse tokens. These tokens serve as input to the WCA module which facilitates the interaction between historical and current tokens. The SiamWCA is elaborated in Sec. 3.4.
Dense feature recovery and detection head. Once the current tokens are augmented with historical information, these sparse tokens are reverted to the x-y plane to form a dense feature map while vacant pillars are filled with zeros. Since LiDAR points only occur on object surfaces, the object centers typically locate at empty space, leading to inaccurate detection. We follow GD-MAE [64] to attach four dense convolutional layers, which spread the feature from occupied pillars to vacant regions. For the detection head, we adopt a center-based head and use the same target assignment strategy as CenterPoint [68].